the point Q is twice as far from (-2,-8) as it is from (13,5). If it lies on the segment joining the two, what are the coordinates of Q?

Answer by stanbon(75887)

the point Q is twice as far from (-2,-8) as it is from (13,5). If it lies on the segment joining the two, what are the coordinates of Q?
---------
Let the coordinates of Q be (x,y)
-----
change in horizontal distance = 13--2 = 15
2x + x = 15
3x = 15
x = 5
2x = 10
-----
The value of "x" is -2+10 = 8
-----
change in vertical distance = 5--8 = 13
2y + y = 13
3y = 13
y = 13/3
2y = 26/3
-------
The value of "y" is -8 + 26/3 = 2/3
====
Ans: Q = (x,y) = (8,2/3)
--------------------------
Cheers,
Stan H.
------------
--